OpenStudy (anonymous):

The sum of two numbers is 30 and their difference is 2. Find the two numbers by writing and solving a system of equations.

OpenStudy (campbell_st):

let the numbers be x and y you can solve this problem by creating 2 equations and solving them simultaneously. x + y = 30 (sum) x - y = 2 (difference) -------- add the 2 equations eliminates y 2x = 32 solve for x. Then substitute the value into either equation to find y. hope it helps